karlieeeeeee:

Is (1, 10) a solution to this system of equations? y= x+5 y= 7x+3

quasia:

Hi~ karlie!! y = 5x 2x + y = 14 Let's take the second equation and isolate y... 2x + y = 14 Let's subtract 2x from both sides... -2x+2x+y=-2x+14 SECOND EQUATION...y=-2x+14 FIRST EQUATION.......y=5x Let's subtract the first equation from the second... 0=-7x+14 Let's add 7x to both sides... 7x+0=7x-7x+14 7x=14 x=2 Let's plug-in x=2 into the first equation.. y=5x y=(5)(2) y=10 Let's plug-in x=2 into the second equation to verify y=10... y=-2x+14 10=[(-2)(2)]+14 10=-4+14 10=10 The two lines meet at (2,10). Your question asked if these lines meet at (10,2) thus rendering such as a "solution". The answer is NO.